y = x³
Find y when x=(3) and (-3)

Answer by stanbon(75887)

y = x^3
Find y when x=(3) and (-3)
-----
If x = 3 then y = 3^3 = 27
-----
If x = -3 then y = (-3)^3 = -27
====================================
Cheers,
Stan H.
=============